OpenStudy (anonymous):

Simplify (x^3/8)^3/4 @Nnesha you know how to do this? If you do can you check it for me

OpenStudy (anonymous):

I got x^9/32 because I believe you just multiply the exponent, which is the fraction

OpenStudy (ybarrap):

Is this the problem? $$ \huge{ \left (\frac{x^3}{8}\right)^{3/4} } $$

OpenStudy (ybarrap):

Or $$ \huge{ \left (x^{3/8}\right)^{3/4} } $$

OpenStudy (anonymous):

the second one, I was trying to draw it but It looked awful...

OpenStudy (ybarrap):

lol, ok

OpenStudy (ybarrap):

Then your answer is correct

OpenStudy (anonymous):

I just multiply the exponents together right(the fractions) which gives me x^9/32?

OpenStudy (anonymous):

Ahh ok thanks I just am checking, If I have any more questions tonight will you be on?

OpenStudy (ybarrap):

yes!

OpenStudy (anonymous):

ok thanks @ybarrap when I do them ill post them and tag you to see if theyre right :)